expand by using binomial theorem: (x^2/2-2/x)^4


1
Expert's answer
2021-08-16T11:15:26-0400

"\left( \cfrac {x^2} 2 - \cfrac 2 x \right) ^{ \large 4} = \left( \cfrac {x^2} 2 \right)^{ \large 4} 
- \large{4} \left(\cfrac {x^2} 2 \right)^{ \large 3}\cfrac 2 x 
+ \large{6} \left( \cfrac {x^2} 2 \right)^{ \large 2}\left( \cfrac 2 x \right)^{ \large 2}
- \large{4}\left( \cfrac {x^2} 2 \right) \left( \cfrac 2 x \right)^{ \large 3} +\\
 +\left( \cfrac 2 x \right)^{ \large 4} = \cfrac {x^8} {16} 
- \large{4} \left(\cfrac {x^6} 8 \right) \cfrac 2 x 
+ \large{6} \left( \cfrac {x^4} 4 \right)\left( \cfrac 4 {x^2} \right)
- \large{4}\left( \cfrac {x^2} 2 \right) \left( \cfrac 8 {x^3} \right) 
 + \cfrac {16} {x^4} = \\
= \cfrac {x^8} {16} - x^5 + \large{6} x^2 - \cfrac {16} x + \cfrac {16} {x^4}"